OpenStudy (anonymous):

. Simplify: 14 - 2 × 6 + 42 22 18 88 42 @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

use PEDMAS first multply: 2*6=?

OpenStudy (anonymous):

12??

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

now subtarct 14-12=?

OpenStudy (anonymous):

2

OpenStudy (alexandervonhumboldt2):

yes now add 2+42=???

OpenStudy (anonymous):

44

OpenStudy (alexandervonhumboldt2):

yes correct, so that is your answer

OpenStudy (anonymous):

but it doesn't have 44 in it can i just select 42?? to see if its correct?

OpenStudy (alexandervonhumboldt2):

42 is not correct

OpenStudy (alexandervonhumboldt2):

some idiot made the question

OpenStudy (anonymous):

lol

OpenStudy (alexandervonhumboldt2):

yay @TheSmartOne this is for you

OpenStudy (anonymous):

Thanks also one more please 5+(x[18/3x2] + 4 @Ale

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (wmj259):

@AlexandervonHumboldt2, hahahahaha

OpenStudy (anonymous):

@wmj259 What??

OpenStudy (wykeise):

heSmartOne

OpenStudy (wykeise):

Thanks also one more please 5+(x[18/3x2] + 4 @Ale

TheSmartOne (thesmartone):

LOL